OpenStudy (anonymous):

Find the equation of the tangent line to the curve at the given point. y = sqrt(x) (9,3). I understand the derivative of f(x) = sqrt(x), I'm just getting lost when trying to work with the points that aren't (1,1)

zepdrix (zepdrix):

So you found the derivative function ok? :) The equation of the line tangent to our function at (9,3) is given by: \[\Large y=mx+b\]Where \(\Large m\), the slope, is \(\Large f'(9)\)

OpenStudy (anonymous):

Yes, the derivative part is easy for me! \[\sqrt{x}\] = 1/2 x. But I guess it's the rest that I'm still not understanding.

zepdrix (zepdrix):

Hmm there should be a square root in your derivative somewhere :D Maybe you just wrote it sloppy. \[\Large f(x)=\sqrt x \qquad\to\qquad f'(x)=\frac{1}{2\sqrt x}\]

zepdrix (zepdrix):

`The slope of the tangent line`, \(\large m\), is given by the derivative function `evaluated at that point`. \[\Large m=f'(9)=\frac{1}{2\sqrt 9}\]

zepdrix (zepdrix):

You're trying to find a line, that `touches` the curve at x=9 and has the same slope. A line is of the form \(\large y=mx+b\). So this problem involves finding 2 unknowns, m and b. m is the tricky part.

zepdrix (zepdrix):

too much? :c confusing? head going to esplode?

OpenStudy (anonymous):

I thiiiiink I got it... haha.

zepdrix (zepdrix):

:3

OpenStudy (anonymous):

Apparently not! I ended up with y-9 = 1/6x (x-3) and then simplified. But that's not the answer. Blaaah.

zepdrix (zepdrix):

Oh you're using point-slope form? ah i see.

zepdrix (zepdrix):

So a line in point slope form:\[\Large y-y_o \quad=\quad m(x-x_o)\]Where the given point was:\[\Large (x_o,\;y_o) \quad=\quad (9,\;3)\]

zepdrix (zepdrix):

So you found m to be,\[\Large f'(9)\quad=\quad \frac{1}{6}\]

zepdrix (zepdrix):

Looks like you were very close. Just one silly little mistake in the middle there :o

OpenStudy (anonymous):

I did it backwards! That's the problem! Yes, 1/6 = m. so the answer is 1/6x + 3/2. Thank you!

zepdrix (zepdrix):

yay good job \c:/

OpenStudy (anonymous):

Thank you very much for your help, a lot more clear now!